OpenStudy (yellowlegoguy99):

Please help with my 8th grade math

OpenStudy (yellowlegoguy99):

Evaluate 27 × a^-2 for a = 3. 1/3 1/9 9 3

OpenStudy (yellowlegoguy99):

I went with D.

OpenStudy (yellowlegoguy99):

So D is right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a^-2= 1/9 27/9=3